OpenStudy (anonymous):

Solve -1/5=6 Option A: -30 Option B: -6/5 Option C: 3/2 Option D: 3

OpenStudy (anonymous):

is that supposed to be -1/5x=6

OpenStudy (anonymous):

yeah aha sorry

OpenStudy (hba):

\[\frac{ -1 }{ 5x }=6\]

OpenStudy (hba):

Its Easy You Can Do It :)

OpenStudy (hba):

Give It A Try :)

OpenStudy (anonymous):

you need to get rid of the -1/5 to get the x alone. in order to do that you have to multiply both sides by the reciprical so (-5/1)(-1/5)x=(-5/1)(6)

OpenStudy (anonymous):

on the left side you are left with 1x and the right side you are left with -30/1 so x=-30

OpenStudy (anonymous):

ty (: